# A Survey on LLM-as-a-Judge

### v5

@@ 4.1. Basic Metric
The main objective of LLM-as-a-judge is to achieve alignment with human judges. Numerous studies approach this by considering the LLM evaluator as a virtual annotator and evaluating the extent of its agreement with human annotators. The percentage agreement metric represents the proportion of samples on which LLM and human annotators agree (Thakur et al., 2024).
@@
where $\mathcal{D}$ is the dataset, $S_{\text{llm}}$ and $S_{\text{human}}$ is the evaluation result of LLM evaluator and human judge respectively, which can be in the form of both score or rank. Additionally, widely used correlation metrics such as Cohen’s Kappa (Thakur et al., 2024) and Spearman’s correlation (Liu et al., 2024b; Bai et al., 2023) are also employed to access agreement. Other works treat the LLM-as-a-judge task as a classification problem, where human annotations serve as the labels, and compute precision, recall, and F1 scores to evaluate the performance (Zhu et al., 2023a; Wang et al., 2023c).
Datasets. Both of above metrics rely on the datasets with LLM-generated response and responding human judgments. Therefore, there is also a practical need to construct a comprehensive benchmark for the meta-evaluation. We list the existing benchmarks and their statistics in Table 1. MTBench (Zheng et al., 2023b) has only 80 human-crafted queries with their responding human annotation and LLMs’ responses. FairEval (Wang et al., 2024c) is constructed from the 80 queries from VicunaBench (Touvron et al., 2023b) with human annotated preference between ChatGPT and Vicuna responses. Chatbot Arena Conversations (Zheng et al., 2023b) is a larger collection of crowdsourced data (about 30k) with human annotated preference. Research (Zeng et al., 2023) construct a benchmark to access the capability of LLM evaluator in evaluating whether a response is following the instruction. This dataset contains human-curated 419 pairs of outputs, one adhering to instructions while the other diverging, yet may possess deceptive qualities that mislead an LLM evaluator. Research (Chen et al., 2024b) evaluate the capabilities of multi-modal LLMs in assisting evaluation tasks across various modalities and introduce MLLM-as-a-Judge, a comprehensive multi-modal benchmark. Recent advances also expand the scope of meta-evaluation benchmarks to specialized domains, including code assessment (Zhao et al., 2024b) and non-English language tasks (Son et al., 2024). Furthermore, CALM (Ye et al., 2024) presents a systematic framework for bias quantification, featuring an automated perturbation mechanism to generate meta-evaluation data for examining 12 distinct types of potential biases in LLM evaluators.
@@
Benchmark Release Year Size Annotation Format Evaluation Dimension Agreement Position Bias Length Bias Bias Types MTBench (Zheng et al., 2023b) 2023 80 Pairwise 3 Chatbot Arena (Zheng et al., 2023b) 2023 30k Pairwise 3 FairEval (Wang et al., 2024c) 2023 80 Pairwise 1 PandaLM (Wang et al., 2023c) 2023 - Pairwise 0 LLMEval2 (Zhang et al., 2023d) 2023 2553 Pairwise 0 Shepherd (Wang et al., 2023b) 2023 1317 Score 0 EvalBiasBench (Park et al., 2024) 2023 80 Pairwise 6 CALM (Ye et al., 2024) 2024 4356 Pairwise & Score 12 JudgeBench (Tan et al., 2024b) 2024 - Pairwise 0 MLLM-as-a-Judge (Chen et al., 2024b) 2024 30k Pairwise & Score 0 CodeJudge (Zhao et al., 2024b) 2024 1860 Score 0 KUDGE (Son et al., 2024) 2024 3324 Pairwise & Score 0
Table 1. Benchmark for meta-evaluation of LLM-judge.
@@ 4.2. Bias
Previous reviews have highlighted that large language models exhibit various types of biases across various tasks (Gallegos et al., 2024a; Dai et al., 2024a; Tan et al., 2024a). These internal biases of LLMs may also affect LLM-as-a-judge, leading to unfair evaluation outcomes and subsequently impacting the development of LLMs. Therefore, it is crucial to understand the types of biases that LLM evaluators might possess and to systematically assess these biases. In this section, we systematically review various types of biases in the LLM-as-a-judge context, including their definitions, relevant metrics, and datasets that can be used for evaluation.
The meta-evaluation of LLM-as-a-judge introduces systematic biases that can be broadly categorized into two classes: task-agnostic biases inherent to LLMs across general applications, and judgment-specific biases unique to LLM-as-a-judge scenarios. This taxonomy aims to clarify their distinct characteristics and implications.
@@ 4.2.1. Task-Agnostic Biases
Self-Enhancement Bias describe the phenomenon that LLM evaluators may prefer response generated by themselves (Zheng et al., 2023b; Ye et al., 2024). This bias has also been known as source bias in retrieval task (Dai et al., 2024b) and open-domain question answering systems (Tan et al., 2024a). Considering the significant self-enhancement bias, as suggested in (Ye et al., 2024), we should avoid using the same model as the evaluator. This is only a stopgap, as we may not use the optimal evaluator when evaluating the most advanced LLMs.
@@ 4.2.2. Judgment-Specific Biases
Judgment-specific biases are either unique to the LLM-as-a-judge setting or have a significant impact on judgment tasks. A classic example is the ”position bias”, which has a more pronounced effect in the context of LLM-as-a-judge where the evaluator often need to compare pairwise responses. Different from task-agnostic biases, judgment-specific biases are more difficult to resolve naturally with the development of foundational large model capabilities and require targeted optimization for judgment tasks.
@@
Position Bias is the tendency of LLM evaluators to favor responses in certain positions within the prompt (Wang et al., 2024c; Thakur et al., 2024; Shi et al., 2024; Ye et al., 2024). This bias may have detrimental effects, as Vicuna-13B could outperform ChatGPT when evaluated by ChatGPT, simply by positioning the response of Vicuna-13B in the second place (Wang et al., 2024c). To measure this bias, recent work (Shi et al., 2024) proposed two metrics: Position Consistency, which quantifies how frequently a judge model selects the same response after changing their positions, and Preference Fairness, which measures the extent to which judge models favor response in certain positions. The study (Wang et al., 2024c) also introduced a metric Conflict Rate to measure the percent of disagreement after change the position of two candidate responses. Their analytical experiments reveal that the degree of positional bias fluctuates depending on the disparity in response quality and the preferred position varies with different LLMs. For instance, GPT-4 tends to favor the first position, while ChatGPT shows a preference for the second position.
@@
Compassion-fade bias describes the effect of the model names (Koo et al., 2023; Ye et al., 2024). This tendency occurs when we explicitly provide model names; for instance, evaluators may be inclined to give higher scores to results labeled as “gpt-4”. This tendency underscores the necessity of anonymous evaluation.
@@
Style Bias refers to the tendency towards certain text style. As revealed in (Chen et al., 2024a), evaluator may also prefer visually appealing content, regardless of its actual validity, such as the text with emoji. Furthermore, LLM evaluators may favor response with certain emotional tones, such as cheerful, sad, angry, and fearful, which is defined as sentiment bias (Ye et al., 2024; Li and Sinnamon, 2023).
@@
Length Bias refers to the tendency to favor responses of a particular length, such as a preference for more verbose responses which is also known as verbosity bias (Park et al., 2024; Huang et al., 2024; Zheng et al., 2023b; Ye et al., 2024). Length bias can be revealed by rephrasing one of the original response into a more verbose one (Ye et al., 2024; Zheng et al., 2023b). Even though these expansions do not introduce new information, there is still concern regarding changes to the original response in terms of perplexity, fluency, or style. Alternatively, previous study (Saito et al., 2023) investigated this bias by comparing multiple sampled responses and revealed a statistical tendency towards longer answers. However, ensuring the comparable quality of multiple samples remains a challenging problem.
@@
Concreteness Bias reflects that LLM evaluators favor responses with specific details, including citation of authoritative sources, numerical values and complex terminologies, which is called authority bias (Park et al., 2024) or citation bias (Chen et al., 2024a; Ye et al., 2024).
The negative effects of concreteness bias arises from the neglect of the factual correctness of these details, thereby encouraging hallucination (Agrawal et al., 2023).
4.2.3. Challenges
To advance the development of LLM-as-a-Judge systems, future efforts should address two key challenges: (i) Need for Systematic Benchmark. Due to the diversity of biases, it is crucial to propose a systematic benchmark to evaluate the extent of various biases. As shown in Table 1, EVALBIASBENCH (Park et al., 2024) was proposed as a test set to measure six types of bias. Other work (Ye et al., 2024) is dedicated to proposing a unified bias testing process, including automated perturbation and a unified metric. They constructed a bias quantification framework CALM, which covers 12 types of bias. Despite these efforts, there is still no systematic benchmark and dataset that includes all types of biases. (ii) Challenges of Controlled Study. When conducting an investigation into a certain type of bias, it is challenging to isolate the specific direction of interest from other biases and quality-related characteristics. For instance, in the case of position bias, lengthening the response could potentially alter the style, fluency, and coherence, or even introduce new biases such as self-enhancement bias. Additionally, the tendency for GPT-4 to favor its own responses over those of GPT-3.5 can be interpreted as either self-enhancement bias or a proper tendency towards higher quality text. Therefore, it is essential for analytical work to carefully control for these variances.
@@ 4.3. Adversarial Robustness
Adversarial robustness refers to the ability of a model to withstand deliberate attempts to manipulate the scores through carefully crafted inputs. Unlike bias evaluations (Section 4.2) which mainly focus on naturally occurring samples, adversarial robustness involves samples intentionally crafted to manipulate scoring, such as inserting phrases that artificially enhance scores. Robustness is crucial because insufficient robustness allows trivial manipulations to deceive the evaluators and to undermine the evaluation of text quality. Ensuring robust evaluators is essential for maintaining accurate and reliable assessments, particularly in high-stakes applications.
Research (Raina et al., 2024) constructed a surrogate model from the black-box LLM-evaluator and the learn a adversarial attack phrases based on it. The evaluation score can be drastically inflated by universally inserting the learned attack phrases without improving the text quality. Similarly, work by Lee et al. (Lee et al., 2024a) introduced EMBER, a benchmark that revealed biases in when assessing outputs with epistemic markers, such as expressions of certainty or uncertainty. Furthermore, other work (Zheng et al., 2024a) demonstrated that even a ”null model” that outputs a constant response irrelevant to input instructions can achieve high win rates for various LLM-as-a-judge methods. Several recent works (Ye et al., 2024; Koo et al., 2023) proposed to increase the evaluation score by adding the majority opinions, such as “90 Other research (Ye et al., 2024; Koo et al., 2023) evaluated robustness against meaningless statement in the System Prompt, e.g., “”Assistant A loves eating pasta”. These works revealed that LLM-as-a-judge are still insufficiently robust against interference irrelevant to text quality. Defensive measures like the perplexity score (Raina et al., 2024; Jain et al., 2023) can only detect limited types of adversarial examples. Therefore, constructing more robust LLM-as-a-judge is a crucial research direction for the future.
5. Meta-evaluation Experiment
@@
In Section 3, we have introduced the improvement strategies adopted by researchers in existing LLM-as-a-judge works to improve the evaluation capabilities of LLM. Although numerous works have proposed meta-evaluation benchmarks to assess the performance of LLMs in evaluation tasks, as shown in Table 1, there is still a lack of meta-evaluation of wh ether these improvement strategies effectively optimize LLM evaluators and which dimensions of evaluation performance are being enhanced. Some improvement strategies may fail to improve the LLM evaluators’ performance or mitigate biases in practical use, leading to computing waste.
In this section, based on the benchmarks mentioned in Section 4, we designed a robust and scalable meta-evaluation tool as shown in Fig. 14, and conducted a simple meta-evaluation experiment on the improvement strategies summarized in Section 3, examining their effectiveness from the perspectives of biases and agreement with human evaluation.
5.1. Experiment Settings
5.1.1. Evaluation Dimensions and Benchmarks
@@ 5.1.2. Evaluation Metrics
For position bias, we use Position Consistency as the metric, which quantifies how frequently the LLM-as-a-judge selects the same response after swapping the position of candidate responses. Formally, given $N$ samples $\{(q_{i},r1_{i},r2_{i})\}^{N}_{i=1}$, for each sample $(q_{i},r1_{i},r2_{i})$, we queried the LLM-as-a-judge with two prompts $P(q_{i},r1_{i},r2_{i})$ and $P(q_{i},r2_{i},r1_{i})$, and obtained two corresponding evaluation results $S_{i}^{r12}$ and $S_{i}^{r21}$. Each $S_{i}$ is $r1_{i}$, $r2_{i}$ or ”TIE”. Then we calculate the position consistency as follows:
Position $\text{Position Consistency}=\frac{\sum_{i=1}^{N}\mathbb{I}(S_{i}^{r12}=S_{i}^{r21})}{N}$
@@
5.1.3. Target LLMs and Strategies
For LLMs, we selected six LLMs commonly used in the automatic evaluation, including closed-source LLMs GPT-4, GPT-3.5, and open-source LLMs Qwen2.5-7B, LLaMA3-8B, Mistral-7B, and Mixtral-8$\times$7B.
@@ 5.1.4. Model Configuration
For open-source LLMs, we adopt Qwen2.5-7B-Instruct [^7] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct, Meta-Llama-3-8B-Instruct [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.3999https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, Mixtral-8$\times$7B-Instruct-v0.1101010https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1, deployed on an Ubuntu machine equipped with a 40GB NVIDIA A100 GPU.
@@
5.2. Experiment Results and Analysis
5.2.1. Comparison with Different LLMs

### v6

@@ 4.1. Agreement with Human Judgments
Given that the initial motivation of LLM-as-a-judge lies in replacing human annotation, the foremost aspect of its evaluation should naturally be the extent to which it aligns with human judgments. Numerous studies approach this by considering the LLM evaluator as a virtual annotator and evaluating the extent of its agreement with human annotators. The percentage agreement metric represents the proportion of samples on which LLM and human annotators agree (Thakur et al., 2024).
@@
where $\mathcal{D}$ is the dataset, $S_{\text{llm}}$ and $S_{\text{human}}$ are the evaluation results of the LLM evaluator and human judge, respectively, which can be in the form of both score or rank. Additionally, widely used correlation metrics such as Cohen’s Kappa (Thakur et al., 2024) and Spearman’s correlation (Liu et al., 2024b; Bai et al., 2023) are also employed to assess agreement. Other works treat the LLM-as-a-judge task as a classification problem, where human annotations serve as the labels, and compute precision and recall to evaluate the performance (Zhu et al., 2023; Wang et al., 2023d).
Benchmark Release Year Size Annotation Format Evaluation Dimension Agreement Position Bias Length Bias Bias Types MTBench (Zheng et al., 2023b) 2023 80 Pairwise 3 Chatbot Arena (Zheng et al., 2023b) 2023 30k Pairwise 3 FairEval (Wang et al., 2024c) 2023 80 Pairwise 1 PandaLM (Wang et al., 2023d) 2023 - Pairwise 0 LLMEval2 (Zhang et al., 2023d) 2023 2553 Pairwise 0 Shepherd (Wang et al., 2023c) 2023 1317 Score 0 EvalBiasBench (Park et al., 2024) 2023 80 Pairwise 6 CALM (Ye et al., 2024) 2024 4356 Pairwise & Score 12 JudgeBench (Tan et al., 2024b) 2024 - Pairwise 0 MLLM-as-a-Judge (Chen et al., 2024b) 2024 30k Pairwise & Score 0 CodeJudge (Zhao et al., 2024b) 2024 1860 Score 0 KUDGE (Son et al., 2024) 2024 3324 Pairwise & Score 0
Table 1. Benchmark for meta-evaluation of LLM-judge.
Datasets. Both of the above metrics rely on the datasets with LLM-generated responses and responding human judgments. Therefore, there is also a practical need to construct a comprehensive benchmark for the meta-evaluation. Table 1 shows existing banchmarks. MTBench (Zheng et al., 2023b) has only 80 human-crafted queries with their corresponding human annotation and LLMs’ responses. FairEval (Wang et al., 2024c) is constructed from the 80 queries from VicunaBench (Touvron et al., 2023b) with human-annotated preference between ChatGPT and Vicuna responses. Chatbot Arena Conversations (Zheng et al., 2023b) is a larger collection of crowdsourced data (about 30k) with human annotated preferences. Research (Zeng et al., 2023) constructs a benchmark to assess the capability of the LLM evaluator in evaluating whether a response follows the instruction. This dataset contains human-curated 419 pairs of outputs, one adhering to instructions while the other diverging, yet may possess deceptive qualities that mislead an LLM evaluator. Research (Chen et al., 2024b) evaluates the capabilities of multi-modal LLMs in assisting evaluation tasks across various modalities and introduces MLLM-as-a-Judge, a comprehensive multi-modal benchmark. Recent advances also expand the scope of meta-evaluation benchmarks to specialized domains, including code assessment (Zhao et al., 2024b) and non-English language tasks (Son et al., 2024). Furthermore, CALM (Ye et al., 2024) presents a systematic framework for bias quantification, featuring an automated perturbation mechanism to generate meta-evaluation data for examining 12 distinct types of potential biases in LLM evaluators.
@@ 4.2. Bias
As LLM-as-a-Judge becomes more widely deployed, it has been observed to manifest a range of salient biases, even in cases where its evaluations align with human judgments. Prior studies have emphasized that large language models inherently exhibit diverse forms of bias across different tasks (Gallegos et al., 2024a; Dai et al., 2024a; Tan et al., 2024a). Such internal biases may propagate into the LLM-as-a-Judge setting, potentially resulting in unfair evaluation outcomes and, in turn, influencing the development of downstream LLMs. Consequently, it is imperative to both characterize the types of biases that LLM evaluators may carry and to establish systematic approaches for their assessment. In this section, we review the major categories of biases in the context of LLM-as-a-Judge, outlining their definitions, associated metrics, and datasets commonly employed for evaluation.
The meta-evaluation of LLM-as-a-judge introduces systematic biases that can be broadly categorized into two classes: task-agnostic biases inherent to LLMs across general applications, and judgment-specific biases unique to LLM-as-a-judge scenarios. This taxonomy aims to clarify their distinct characteristics and implications. The former may be partially attributed to the characteristics of the judgment task itself, such as its specific input-output format, which could potentially be mitigated through task-specific design. In contrast, task-agnostic biases are more fundamental issues inherent to the LLMs themselves and are therefore more difficult to address. Mitigating such biases likely depends on advancements in the foundation models.
@@ 4.2.1. Task-Agnostic Biases
Self-Enhancement Bias describes the phenomenon that LLM evaluators may prefer responses generated by themselves (Zheng et al., 2023b; Ye et al., 2024). This bias has also been known as source bias in retrieval task (Dai et al., 2024b) and open-domain question answering systems (Tan et al., 2024a). Considering the significant self-enhancement bias, as suggested in (Ye et al., 2024), we should avoid using the same model as the evaluator. This is only a stopgap, as we may not use the optimal evaluator when evaluating the most advanced LLMs.
@@ 4.2.2. Judgment-Specific Biases
Judgment-specific biases are either unique to the LLM-as-a-judge setting or have a significant impact on judgment tasks. A classic example is the ”position bias”, which has a more pronounced effect in the context of LLM-as-a-judge, where the evaluator often needs to compare pairwise responses. Different from task-agnostic biases, judgment-specific biases are more difficult to resolve naturally with the development of foundational large model capabilities and require targeted optimization for judgment tasks.
@@
Position Bias is the tendency of LLM evaluators to favor responses in certain positions within the prompt (Wang et al., 2024c; Thakur et al., 2024; Shi et al., 2024; Ye et al., 2024). This bias may have detrimental effects, as Vicuna-13B could outperform ChatGPT when evaluated by ChatGPT, simply by positioning the response of Vicuna-13B in the second place (Wang et al., 2024c). To measure this bias, recent work (Shi et al., 2024) proposed two metrics: Position Consistency, which quantifies how frequently a judge model selects the same response after changing their positions, and Preference Fairness, which measures the extent to which judge models favor a response in certain positions. The study (Wang et al., 2024c) also introduced a metric Conflict Rate to measure the percent of disagreement after changing the position of two candidate responses. Their analytical experiments reveal that the degree of positional bias fluctuates depending on the disparity in response quality, and the preferred position varies with different LLMs. For instance, GPT-4 tends to favor the first position, while ChatGPT shows a preference for the second position.
@@
Compassion-fade bias describes the effect of the model names (Koo et al., 2023; Ye et al., 2024). This tendency occurs when we explicitly provide model names. for instance, evaluators may be inclined to give higher scores to results labeled as “gpt-4”. This tendency underscores the necessity of anonymous evaluation.
@@
Style Bias refers to the tendency towards a certain text style. As revealed in (Chen et al., 2024a), an evaluator may also prefer visually appealing content, regardless of its actual validity, such as the text with emojis. Furthermore, LLM evaluators may favor responses with certain emotional tones, such as cheerful, sad, angry, and fearful, which is defined as sentiment bias (Ye et al., 2024; Li and Sinnamon, 2023).
@@
Length Bias refers to the tendency to favor responses of a particular length, such as a preference for more verbose responses, which is also known as verbosity bias (Park et al., 2024; Huang et al., 2024; Zheng et al., 2023b; Ye et al., 2024). Length bias can be revealed by rephrasing one of the original responses into a more verbose one (Ye et al., 2024; Zheng et al., 2023b). Even though these expansions do not introduce new information, there is still concern regarding changes to the original response in terms of perplexity, fluency, or style. Alternatively, a previous study (Saito et al., 2023) investigated this bias by comparing multiple sampled responses and revealed a statistical tendency towards longer answers. However, ensuring the comparable quality of multiple samples remains a challenging problem.
@@
Concreteness Bias reflects that LLM evaluators favor responses with specific details, including citation of authoritative sources, numerical values, and complex terminologies, which is called authority bias (Park et al., 2024) or citation bias (Chen et al., 2024a; Ye et al., 2024). The negative effects of concreteness bias arise from the neglect of the factual correctness of these details, thereby encouraging hallucination (Agrawal et al., 2023).
@@ 4.3. Adversarial Robustness
As LLM-as-a-Judge becomes further integrated into evaluation pipelines and increasingly serves as the standard protocol, a variety of adversarial attacks have emerged. Adversarial robustness refers to the ability of a model to withstand deliberate attempts to manipulate the scores through carefully crafted inputs. Unlike bias evaluations (Section 4.2), which mainly focus on naturally occurring samples, adversarial robustness involves samples intentionally crafted to manipulate scoring, such as inserting phrases that artificially enhance scores. Robustness is crucial because insufficient robustness allows trivial manipulations to deceive the evaluators and to undermine the evaluation of text quality. Ensuring robust evaluators is essential for maintaining accurate and reliable assessments, particularly in high-stakes applications.
Research (Raina et al., 2024) constructed a surrogate model from the black-box LLM-evaluator and then learn a adversarial attack phrases based on it. The evaluation score can be drastically inflated by universally inserting the learned attack phrases without improving the text quality. Similarly, work by Lee et al. (Lee et al., 2024a) introduced EMBER, a benchmark that revealed biases in when assessing outputs with epistemic markers, such as expressions of certainty or uncertainty. Furthermore, other work (Zheng et al., 2024a) demonstrated that even a ”null model” that outputs a constant response irrelevant to input instructions can achieve high win rates for various LLM-as-a-judge methods. Similarly, recent work (Zhao et al., 2025) also revealed that on-word symbols (e.g., “:”) or reasoning openers like “Thought process:” can often fool LLM evaluators to produce positive evaluations. Several recent works (Ye et al., 2024; Koo et al., 2023) proposed to increase the evaluation score by adding the majority opinions, such as “90% believe this is better”. Other research (Ye et al., 2024; Koo et al., 2023) evaluated robustness against meaningless statement in the System Prompt, e.g., ”Assistant A loves eating pasta”. These works revealed that LLM-as-a-judge are still insufficiently robust against interference irrelevant to text quality. Defensive measures like the perplexity score (Raina et al., 2024; Jain et al., 2023) can only detect limited types of adversarial examples. Therefore, constructing more robust LLM-as-a-judge is a crucial research direction for the future.
4.4. Empirical Experiment
@@
In Section 3, we have introduced the improvement strategies in existing LLM-as-a-judge works to improve the evaluation capabilities of LLMs. Although numerous works have proposed meta-evaluation benchmarks to assess the performance of LLMs in evaluation tasks, there is a lack of meta-evaluation on whether these improvement strategies effectively make LLM-as-a-Judge more reliable. So in the preceding subsection of this Section 4, we systematically introduced the dimensions and criteria for evaluating the evaluation performance and reliability of LLM-as-a-Judge, thereby providing quantifiable metrics that inform the evaluation of the aforementioned strategies.
Based on partial dimensions mentioned in sec 4.1 and sec 4.2, we designed a robust and scalable meta-evaluation tool as shown in Figure 13 and conducted a simple meta-evaluation experiment on the improvement strategies summarized in Section 3, examining their effectiveness from the perspectives of biases and agreement with human evaluation.
4.4.1. Experiment Settings
Evaluation Dimensions and Benchmarks
@@ 4.4.1. Experiment Settings
For position bias, we use Position Consistency as the metric, which quantifies how frequently the LLM-as-a-judge selects the same response after swapping the position of candidate responses. Formally, given $N$ samples $\{(q_{i},r1_{i},r2_{i})\}^{N}_{i=1}$, for each sample $(q_{i},r1_{i},r2_{i})$, we queried the LLM-as-a-judge with two prompts $P(q_{i},r1_{i},r2_{i})$ and $P(q_{i},r2_{i},r1_{i})$, and obtained two evaluation results $S_{i}^{r12}$ and $S_{i}^{r21}$. Each $S_{i}$ is $r1_{i}$, $r2_{i}$ or ”TIE”. Then we calculate the Position Consistency as follows:
$\text{Position Consistency}=\frac{\sum_{i=1}^{N}\mathbb{I}(S_{i}^{r12}=S_{i}^{r21})}{N}$
@@
Target LLMs and Strategies
[…]
@@
[…]
@@
4.4.2. Results and Analysis
Comparison with Different LLMs